What are the nominal and effective costs of trade credit under the credit terms of 3/20, net 40? Assume a 365-day year. Do not round intermediate calculations. Round your answers to two decimal places. Nominal cost of trade credit: % Effective cost of trade credit: %.
Answer #1

Nominal cost of trade credit = (3/97) × [365 / (40 - 20)] = 0.5644 or 56.44%

Effective cost of trade credit = [1 + (3/97)][365 / (40 - 20)] - 1 = 0.7435 or 74.35%